Hi. How do you show the steps and solve this:
4X+2(x-1)= -20

Answer by stanbon(75887)

4X+2(x-1)= -20
4x + 2x-2 = -20
6x -2 = -20
6x = -18
x = -3
--------------
Cheers,
Stan H.